Please help me :(
6(x+2)+4x=5x-23
Thank you.

Answer by uma(370)

6(x+2) + 4x = 5x - 23
==> 6x + 12 + 4x = 5x - 23 [removing the parenthesis]
==> 10x + 12 = 5x - 23
==> 10x + 12 - 5x = 5x - 23 - 5x [subtracting 5x from both the sides]
==> 5x + 12 = - 23
==> 5x + 12 - 12 = - 23 - 12
==> 5x = - 35
==-> 5x/5 = -35/5
==> x = -7

Good Luck!!!